OpenStudy (anonymous):

Medal to whoever can help me solve! The function f is defined by f(x)=x^2+2 Find f(5n).

OpenStudy (anonymous):

replace x by 5n

OpenStudy (muzzack):

well replace f(x) with 5n and eith all x's

OpenStudy (anonymous):

Yeah, I did. And when I went to solve it I got slightly confused, I haven't used algebra since like last year, and unfortunately its kinda kicking my retrice

OpenStudy (anonymous):

Because then you get 5n^2+2

OpenStudy (anonymous):

\[(ab)^2 = a^2b^2\]

OpenStudy (anonymous):

Replacing it with 5n would yield answer 25n^2 + 2

OpenStudy (anonymous):

you have to square the 5 as well

OpenStudy (anonymous):

how do you get rid of the n?

OpenStudy (anonymous):

n is your friend. no need to get rid of it

OpenStudy (anonymous):

so 25n^2+2

OpenStudy (muzzack):

you have to isolate it

OpenStudy (anonymous):

okay, so would you do the square root to get the square to disappear

OpenStudy (anonymous):

no. that is your answer. f(5n) = 25n^2 + 2

OpenStudy (anonymous):

n is just a variable

OpenStudy (anonymous):

oh.. well I just made myself look really stupid. lol

OpenStudy (anonymous):

f(3) would be (3)^2 + 2 = 9 + 2 = 11 n is just undefined

OpenStudy (anonymous):

no you didn't. this is a place to learn ^_^

OpenStudy (muzzack):

yes @Euler271

OpenStudy (anonymous):

Okay, well thank you everybody! You all were very helpful!